Mr. Mendeleev found elements known as element X and element Y, he found that element X and element Y are one of the transition element and non-metallic respectively in periodic Table


Identify Element X and Element Y to their groups from Periodic Table.


Using the information from Table 1


i) calculate the mass of atom X assming the number of mole similar with Zinc


ii) calculate the number of atom Y assming the mass of element Y similar with 16. 3 g of sulphur



Table 1


Elements : zinc

Number of moles : 0.365

Mass, g : 23.3

Number of atoms : __


Elements : Sulphur

Number of moles : __

Mass, g : 16.3

Number of atoms : 3.06x10^23


1
Expert's answer
2021-04-21T04:03:53-0400

Table 1: Number of atoms: 0.365*6.02*1023 = 2.197*1023


Table 2: Number of moles: 3.06/6.02 = 0.508